OpenStudy (dabbingpikachu):

Part B: If (60)x = 1, what are the possible values of x? Explain your answer. (5 points)

OpenStudy (dabbingpikachu):

Part B: If (6^0)x = 1, what are the possible values of x? Explain your answer. (5 points) *Correction*

OpenStudy (leenathan):

il that username XD

OpenStudy (dabbingpikachu):

Thank you :D

OpenStudy (leenathan):

ok s0 6 to the power of 0 = 1 so x can be 0

OpenStudy (leenathan):

\[6^0=1\]

OpenStudy (leenathan):

\[(6^0)x = 1 x ...can..be ..0\]

OpenStudy (kayleerose0715):

its going to be 1/60 becasue your going to divide by 60 on both sides to make it equal so it will look like (60)1/60=1 and thats true 1/60 times 60 equals 1

OpenStudy (leenathan):

ur wrong

OpenStudy (leenathan):

thats not the one were doing

OpenStudy (dabbingpikachu):

IDK WHO TO BELIEVE

OpenStudy (leenathan):

Part B: If (\[6^0\]x = 1, what are the possible values of x? Explain your answer. (5 points) *Correction*

OpenStudy (kayleerose0715):

how its a true statement and it says there can be multiple answers

OpenStudy (leenathan):

a number to the power of 0 = 1

OpenStudy (leenathan):

you doing the wrong qustion

OpenStudy (kayleerose0715):

yeah so 60 *1 =60 not 1 its supposed to equal 1

OpenStudy (leenathan):

\[6^0 *x=1\]

OpenStudy (leenathan):

its not 60

OpenStudy (leenathan):

its \[6^0\]

OpenStudy (leenathan):

that why it has ^ ....6^0

OpenStudy (dabbingpikachu):

SOMEONE JUST SAY WHOSE CORRECT

OpenStudy (leenathan):

me

zepdrix (zepdrix):

Are you sure the original equation wasn't meant to be this?\[\large\rm (6^0)^x=1\]Just checking.

OpenStudy (leenathan):

im a human calculator and i have 112 fan i obviously am right

OpenStudy (dabbingpikachu):

Yes zepdrix

OpenStudy (leenathan):

then im right

OpenStudy (kayleerose0715):

so then yeah he would be correct i was solving the original one

OpenStudy (leenathan):

medal plz

OpenStudy (dabbingpikachu):

Bow

OpenStudy (leenathan):

i now have 216 medals!!!!!!!!!!!!!!!

OpenStudy (leenathan):

like 2016

OpenStudy (dabbingpikachu):

Well now this is full of a bunch of stuff that i have to dig through to find the original answer

OpenStudy (leenathan):

its 0

OpenStudy (leenathan):

x=0

OpenStudy (kayleerose0715):

and btw the followers dont mean that your smarter im doing college algebra 2 years ahead of what my class actually is supposed to do this year please do not use followers as an excuse

OpenStudy (leenathan):

\[6^0 * x =1 \] x = 0

OpenStudy (leenathan):

oh nice but im mostly correct

OpenStudy (kayleerose0715):

yes i didnt realize there was a correction i thought there were two different questions to answer i was answering the first one before i got to the second one